How do you write the expression (-6-29i)+(8-12i) as a complex number in standard form?
Answer by Earlsdon(6294)

The standard form for complex numbers is: (a+bi), so...
(-6-29i)+(8-12i) add the real parts and add the imaginary parts.
-6+8 = 2
-29i+(-12i) = -41i Put them together to get:
(2-41i)